### v1

# Four-Plane Factorized Video Autoencoders

Abstract
Latent variable generative models have emerged as powerful tools for generative tasks including image and video synthesis. These models are enabled by pretrained autoencoders that map high resolution data into a compressed lower dimensional latent space, where the generative models can subsequently be developed while requiring fewer computational resources. Despite their effectiveness, the direct application of latent variable models to higher dimensional domains such as videos continues to pose challenges for efficient training and inference. In this paper, we propose an autoencoder that projects volumetric data onto a four-plane factorized latent space that grows sublinearly with the input size, making it ideal for higher dimensional data like videos. The design of our factorized model supports straightforward adoption in a number of conditional generation tasks with latent diffusion models (LDMs), such as class-conditional generation, frame prediction, and video interpolation. Our results show that the proposed four-plane latent space retains a rich representation needed for high-fidelity reconstructions despite the heavy compression, while simultaneously enabling LDMs to operate with significant improvements in speed and memory.
1 Introduction
A defining trait of recent advances in image and video generation is that, as models grow more powerful, they increasingly push against the boundaries of current computational limits. Despite their impressive generative capabilities, these models’ vast resource demands hinder scalability and discourage widespread deployment. Naturally, improving the efficiency of these generative models has become an active research concern [20, 56, 55, 1].
One effective strategy to make generative modeling computationally feasible is through latent modeling, which enables generative models to operate on lower-dimensional representations [3, 38, 5, 34, 11, 20, 35, 18, 1]. Latent models use autoencoders to compress high-resolution visual data into a more compact latent space, where the generative process can proceed with a reduced computational burden – the generative model’s memory footprint and training and inference times are all directly improved by working in the smaller latent space.
Figure 1: Factorized latent representation. Traditional volumetric latents in diffusion models yield a sequence length of $t\times h\times w$ (top row), which scales linearly with the input size and demands high computational resources. Our proposed factorized representation reduces sequence length to $t\times(h+w)+2\times h\times w$ (bottom row), achieving a more compact latent space that scales sublinearly with input size, enabling faster, more efficient video generation; while maintaining high quality.
However, in typical autoencoders, the resulting latent size still scales linearly with the original input size, so the compression offers only a limited benefit when deployed in high-dimensional domains, such as handling videos with high spatial and temporal resolutions [20, 35] (Figure 1).
In this paper, we explore improving the efficiency of latent generative models through more aggressive reduction of the latent resolution. The central objective is to achieve this compression without sacrificing representation quality. To address this challenge, we propose a novel four-plane factorized latent autoencoder that maps volumetric space-time signals onto a latent space through four axis-aligned planar projections. Since the orthogonal projections capture complementary features of the space-time volume, the original signal can still be reconstructed from this more compact latent space with high fidelity. We summarize the key attributes of our contribution below:
  • The four-plane factorized autoencoder offers a significant reduction in latent resolution, as its size scales sublinearly with the total input size (see Figure 1). At the same time, signals decoded from the factorized latents are of comparable fidelity to those reconstructed from larger volumetric latents.
  • Our factorized autoencoder presents an effective latent representation for generative tasks – for example a typical transformer-based diffusion model used on volumetric features [20] generates samples twice as fast when trained on our factorized space, while maintaining comparable generation quality.
  • The latent structure is easily adaptable to a variety of image-conditioned generation tasks. We demonstrate this property on two common video applications, two-frame interpolation and frame extrapolation.
Across a variety of tasks, our experiments suggest the proposed four-plane factorized autoencoder provides an efficient alternative for generative models that traditionally operate on volumetric latent spaces.
2 Related work
2.1 Diffusion models for video synthesis
Denoising Diffusion Probabilistic Models (DDPM) [22] introduced a novel method for generating images by iteratively denoising a sequence of noisy images. This approach has been highly successful for both image [13, 40, 34, 11, 25] and video synthesis [24, 4, 23, 49, 19, 3].
Of the more recent diffusion models developed for video generation, many operate on a volumetric spatiotemporal latent space. VDM [24] employs a 3D U-Net autoencoder architecture [10, 39] to learn this space. The factorized space-time attention mechanism enables joint training on both images and videos. This architecture illustrates that handling spatial and temporal information simultaneously leads to more coherent video sequences.
To address scalability for high resolution video generation, Imagen Video [23] extends VDM by introducing a cascade of models that essentially alternate temporal and spatial superresolution. Lumiere [2] introduced the STUNet architecture, which generates entire videos directly with improved temporal coherence. VideoLDM [4] constructs a video model starting with pretrained image models and inserting temporal layers before fine-tuning on high-quality videos.
2.2 Video tokenizers
Many of the latent video diffusion models highlighted above rely on some form of video tokenization to compress high dimensional videos into a compact latent space. The pioneering vector quantization approaches for image tokenization, for example VQ-VAE [46], VQ-VAE2 [36], and VQGAN [15], can be applied to videos in a frame-by-frame manner. MAGVIT [51] introduced a 3D-VQ autoencoder to quantize video data into spatio-temporal tokens, making it a powerful tool for a range of video generation tasks such a frame prediction, video inpainting, and frame interpolation. MAGVIT-v2 [52] introduced significant advancements, including a lookup-free quantization method, which allows for an expanded vocabulary without compromising performance. Additionally, MAGVIT-v2 enables joint image and video modeling through a causal 3D CNN architecture. The MAGVIT-v2 tokenizer was used successfully in W.A.L.T [20] for photorealistic image and video generation. The early 3D encoder layers in our factorized architecture are based on MAGVIT-v2.
2.3 Video frame interpolation
Video frame interpolation [30, 14] has distinct interpretations depending on the temporal distance between frames. In the case where there is significant motion between frames, the challenging task can be addressed by generative models. The application of our factorized latent representation to frame interpolation can be categorized as a two-frame conditioned diffusion model. The first such effort to use a latent diffusion model was LDMVFI [12]. In contrast, ViDiM [28] models in pixel space and generates the entire video at once improving temporal consistency. To improve interpolation quality, ViDiM employs a cascaded diffusion approach: it first generates a low-resolution video, followed by an upsampling diffusion model that refines the output to higher resolutions
2.4 Tri-plane factorization
Tri-plane representations, which factorize volumetric data into three orthogonal 2D planes, have been widely used as compact representations of 3D neural fields [42, 7, 16]. Coupled with diffusion model architectures suited to planar representations, they have been used for a variety of applications, such as textured 3D model generation [50], 3D neural field generation [42], and semantic scene generation [31]. The same concept was applied to videos in PVDM [54], where encoding videos to tri-plane latent features enables the use of a 2D U-Net architecture [13] for training diffusion models, bringing significant improvements in efficiency. In contrast, our work proposes a distinctive four-plane factorization approach, which broadens the model’s scope, enabling tasks like frame extrapolation and video interpolation.
Figure 2: Model overview. The tokenizer first maps the input video into a volumetric latent representation through $3D$ convolutional architecture, which is then factorized into four planes. Temporal planes are created by mean pooling along the height and width dimensions, capturing time-varying features. Spatial planes are obtained by splitting along the time axis and independently averaging along this dimension, focusing on spatial consistency (highlighted in green). During decoding, the four planes are mapped back into a volume: for each spatial-temporal location, features from the corresponding four planes (highlighted in blue) are concatenated to reconstruct the full volumetric features. These combined features are passed through a decoder to produce the final output video.
3 Factorized video latent representations
In latent video diffusion models, a key component is the autoencoder, which compresses the input video data into a compact latent representation. To achieve this, prior works [20, 21] employ a 3D convolutional architecture that encodes the 3D video volume $\mathbf{X}\in\mathbb{R}^{T\times H\times W\times C}$ into a feature volume $\mathbf{Z}\in\mathbb{R}^{t\times h\times w\times c}$, where $t=\frac{T}{f_{t}}$, $h=\frac{H}{f_{s}}$, $w=\frac{W}{f_{s}}$. Here $f_{t}$ and $f_{s}$ are the temporal and spatial downsampling factors. The channel dimension $c$ is typically expanded (e.g., $c=8$ is a common choice of channel dimension that balances autoencoder reconstruction and diffusion performance).
While this compression does offer significant reduction in the spatial and temporal resolution, the total size ($t\times h\times w$) still scales linearly with the input size. For computationally expensive generative models such as transformer [47]-based diffusion models, this sequence length would still pose a challenge. Improving the efficiency of transformer-based models can either be achieved by addressing the design of the model itself (e.g., sub-quadratic attention mechanisms). or by decreasing the sequence length. In this work we explore the latter by introducing a four-plane factorized autoencoder that we describe in the following section.
3.1 Four plane factorization
Our factorization approach is intended to directly and simply reduce the cubic footprint of volumetric spatiotemporal latent spaces. At a high level, our approach decomposes the 3D feature volume into four distinct planes, each capturing complementary aspects of the video’s spatial and temporal features. This leads to a more efficient representation that accelerates training while preserving the necessary information for high-quality reconstruction. Our introduction of a four-plane factorization is motivated by the versatility it brings, making it adaptable to a range of tasks such as class-conditional generation, frame extrapolation, and video interpolation.
3.1.1 Factorization.
Given an input video $\mathbf{X}\in\mathbb{R}^{T\times H\times W\times 3}$, our encoder network first converts it into a feature volume $\mathbf{Z}\in\mathbb{R}^{t\times h\times w\times c}$ using a causal 3D convolution architecture similar to the one introduced in MAGVIT-v2 [53]. The feature volume is then factorized into four planes along three directions: two spatial planes, $\mathbf{P}_{xy}^{1},\mathbf{P}_{xy}^{2}\in\mathbb{R}^{h\times w\times c}$, aligned along the $xy$-dimension, and two spatiotemporal planes, $\mathbf{P}_{xt}\in\mathbb{R}^{t\times h\times c}$ and $\mathbf{P}_{yt}\in\mathbb{R}^{t\times w\times c}$, aligned along the $xt$ and $yt$ dimensions, respectively.
To obtain the latent feature planes, we apply a factorization operation along certain dimensions. Specifically, the two spatio-temporal planes $\mathbf{P}_{xt}$ and $\mathbf{P}_{yt}$ are obtained by applying a factorization operation along the height and width dimensions, respectively:
$\mathbf{P}_{xt}$ $=\Lambda_{w}(\mathbf{Z})\in\mathbb{R}^{t\times h\times c}$ (1)
$\mathbf{P}_{yt}$ $=\Lambda_{h}(\mathbf{Z})\in\mathbb{R}^{t\times w\times c}$ (2)
where $\Lambda_{h}$ and $\Lambda_{w}$ performs pooling or dimensionality reduction across spatial dimensions $h$ and $w$ resulting in the planar representation.
Similarly, the spatial planes contain the temporally aggregated features across frames, capturing the spatial structure and background information in the video. To obtain the spatial planes $\mathbf{P}_{xy}^{1}$ and $\mathbf{P}_{xy}^{2}$, we adapt our factorization approach based on the application. For class-conditional generation and frame prediction, we split the latent feature volume $\mathbf{Z}\in\mathbb{R}^{t\times h\times w\times c}$ into two segments along the temporal dimension, then aggregate each segment over $T$, yielding:
$\mathbf{P}_{xy}^{1}=\Lambda_{t}(\mathbf{Z}_{1:\lfloor\frac{t}{2}\rfloor}),\quad\mathbf{P}_{xy}^{2}=\Lambda_{t}(\mathbf{Z}_{\lceil\frac{t}{2}\rceil:t}),$ (3)
[…]
(UCF dataset) (Kinetics-600 dataset)
Video Diffusion [24] - 16.2 1.1B 256
RIN [27] - 10.7 411M 1000
TATS [17] 332 - 321M 1024
Phenaki [48] - 36.4 227M 48
MAGVIT [51] 76 9.9 306M 12
MAGVIT-v2 [53] 58 4.3 307M 24
W.A.L.T [20] 46 TML]B3E5FC3.3 313M 50
W.A.L.T* [20] TML]E3F6FF39 TML]E3F6FF5.7 313M 50
Ours TML]B3E5FC38 8.6 214M 50
Table 1: Class-conditional generation on UCF and frame prediction on Kinetics-600. WALT* represents our re-training and re-evaluation of the WALT baseline. Our method achieves competitive performance with WALT on the UCF task and performs slightly lower on the K600 frame prediction task, showcasing efficient performance across both datasets.
For the interpolation task, where the first and last frames are available, we leverage this information by using a pre-trained image tokenizer to directly obtain the latent representations of these frames. Specifically, we set:
$\mathbf{P}_{xy}^{1}=E(\mathbf{X}_{0}),\quad\mathbf{P}_{xy}^{2}=E(\mathbf{X}_{T}),$ (4)
where $E$ denotes the pretrained image tokenizer and $\mathbf{X}_{0}$ and $\mathbf{X}_{T}$ are the boundary frames. This approach effectively incorporates key frame information into the spatial planes, enhancing the model’s interpolation accuracy.
3.1.2 Recomposition
Given the four latent planes $\mathbf{P}_{xy}^{1},\mathbf{P}_{xy}^{2},\mathbf{P}_{xt},$ and $\mathbf{P}_{yt}$, the decoder reconstructs the input video by first reconstituting these planes back into a 3D feature volume. To utilize existing 3D convolutional architectures, we construct an intermediate volume $\mathbf{V}\in\mathbb{R}^{t\times h\times w\times c}$ by back-projecting features from each plane onto corresponding locations within the target volume dimensions $(t,h,w)$.
For any spatial-temporal location $(x,y,t)$ in the volume, we extract features from each of the planes by projecting onto their respective dimensions (depicted in the blue box in Figure 2). Specifically:
$\mathbf{f}_{xy}^{1}=\mathbf{P}_{xy}^{1}(x,y),$ $\quad\mathbf{f}_{xy}^{2}=\mathbf{P}_{xy}^{2}(x,y)$
$\mathbf{f}_{xt}=\mathbf{P}_{xt}(x,t),$ $\quad\mathbf{f}_{yt}=\mathbf{P}_{yt}(y,t).$
Here $\mathbf{f}_{xy}^{1}$, $\mathbf{f}_{xy}^{2}$, $\mathbf{f}_{xt}$, and $\mathbf{f}_{yt}$ will contain features queried from their respective planes, using the corresponding spatial or temporal coordinates. These features are then combined using an operation such as element-wise addition or concatenation, yielding:
$\mathbf{V}(x,y,t)=\mathrm{Combine}(\mathbf{f}^{1}_{xy},\mathbf{f}^{2}_{xy},\mathbf{f}_{xt},\mathbf{f}_{yt}),$
where $\mathrm{Combine}$ represents the chosen operation, e.g., $\mathrm{Combine}(\cdot)=\sum$ for summation, or $\mathrm{Combine}(\cdot)=\mathrm{concat}(\cdot)$ for concatenation, along the channel dimension.
This reconstructed feature volume $\mathbf{V}$ is then fed into a decoder with a structure similar to MAGVIT-v2, which progressively upsamples the features and applies 3D convolutions to generate the final video $\hat{\mathbf{X}}\in\mathbb{R}^{T\times H\times W\times 3}$.
3.2 Generative modeling with factorized latents
With a trained factorized latent model, we obtain a compact, efficient representation of input video data, suitable for training generative models. In our experiments we utilize established techniques for transformer-based latent diffusion models.
Latent diffusion models (LDMs) gradually transform the latent representation of data into noise in a forward diffusion process, then reverse this transformation to generate new samples. Given an initial latent sample $\mathbf{x}_{0}$, the forward process generates a sequence of increasingly noisy latents $\{\mathbf{z}_{t}\}_{t=1}^{T}$ as follows:
$\mathbf{z}_{t}=\sqrt{\alpha_{t}}\mathbf{z}_{t-1}+\sqrt{1-\alpha_{t}}\mathbf{\epsilon},\quad\mathbf{x}_{0}\approx D(\mathbf{z}_{0}),$ (5)
where $\alpha_{t}$ controls the noise schedule and $D$ decodes the final latent back to data space. The reverse process, defined by $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t})=\mathcal{N}(\mathbf{z}_{t-1};\mu_{\theta}(\mathbf{z}_{t},t),\sigma_{t}^{2}\mathbf{I})$, seeks to denoise the latent variables and reconstruct the original data distribution. This denoising is learned by minimizing a variational lower bound, often simplified into practical objectives [22]. Here, we adopt the v-parameterization, following recent diffusion improvements [41].
To train a LDM on our factorized representation we use a transformer architecture similar to W.A.L.T. [20]. We create a 1D sequence by flattening the four planes—$\mathbf{P}_{xt}$, $\mathbf{P}_{yt}$, $\mathbf{P}^{1}_{xy}$, and $\mathbf{P}^{2}_{xy}$—and concatenating them along the sequence length dimension. This results in a sequence length of $h\times t+w\times t+2\times h\times w$ (as shown in Figure 1).
4 Experiments
In order to evaluate the performance of the proposed four-plane factorized autoencoder, we combine it with a diffusion model (Section 3.2) and apply it towards three different experimental benchmarks: class-conditional video generation, future frames prediction, and two-frame interpolation. The observed performance on these diverse synthesis experiments demonstrates the versatility of the proposed four-plane latent space.
Figure 3: Class-conditional generation results on the UCF dataset. The model generates 17 frames at a resolution of 256 x 256. Here, we display 9 frames, selected by subsampling every other frame (stride of 2), from W.A.L.T and our model, to illustrate the continuity and quality of generated video sequences.
4.1 Class-conditional generation and future frames prediction
We introduce the details of the class-conditional generation and frame prediction tasks together since their experimental settings share some similarities.
Datasets.
The class-conditional generation experiment is performed on the UCF-101 dataset [44] which contains 9,537 videos across 101 categories, covering a wide range of actions in total. The future frames prediction experiment is performed on the Kinetics-600 [6] dataset, consisting of a nearly $400,000$ video clips sourced from YouTube, containing around 600 action classes that cover a wide range of human activities.
Autoencoder pre-training.
Following W.A.L.T. [20], we pre-train the factorized autoencoder on the Kinetics-600 dataset for both the class-conditional generation and frame extrapolation tasks.
Diffusion model training.
For the class-conditional generation task, we train the diffusion model to generate four latent feature planes that decode to $128\times 128$ resolution videos with 17 frames, conditioned on the class label. This setup aligns with the approach used in W.A.L.T. [20] to ensure a fair comparison.
For the frame prediction task, we condition on the first spatial plane, $P_{xy}^{1}$, and generate the remaining three planes. Leveraging a causal encoder, as in MAGVIT-v2, enables this conditional generation to parallel the setup in W.A.L.T.’s frame prediction experiments, where the model is conditioned on two latent frames.
4.1.1 Metric
To evaluate the quality of generated videos, we use the Fréchet Video Distance (FVD) [45] as our primary metric. FVD measures the similarity between the distributions of generated and real videos, assessing both spatial realism and temporal coherence. Formally, FVD is defined as:
$\text{FVD}=\|\mu_{\text{real}}-\mu_{\text{gen}}\|^{2}+\operatorname{Tr}(\Sigma_{\text{real}}+\Sigma_{\text{gen}}-2(\Sigma_{\text{real}}\Sigma_{\text{gen}})^{1/2}),$
where $\mu_{\text{real}},\Sigma_{\text{real}}$ and $\mu_{\text{gen}},\Sigma_{\text{gen}}$ are the mean and covariance matrices of feature embeddings for real and generated
Davis-7 UCF-7
[…]
Flow Based AMT [32] 21.09 0.544 0.254 234.5 26.06 0.813 0.144 344.5
RIFE [26] 20.48 0.511 0.258 240.0 25.73 0.804 0.135 323.8
FILM [37] 20.71 0.528 0.270 214.8 25.90 0.811 0.137 328.2
[…]
VIDIM [29] 19.62 0.470 0.257 199.3 24.07 0.781 0.149 278.0
Ours 19.47 0.446 0.256 156.1 24.00 0.769 0.141 216.9
Table 2: Video interpolation results on DAVIS-7 and UCF-7. Our method is compared against several video interpolation baselines, assessing both reconstruction and generative metrics, across all 7 interpolated frames. See the text for additional discussion.
4.1.2 Analysis
As shown in  Table 1, our diffusion model, trained on the proposed factorized latent representation, outperforms most prior works and performs comparably to W.A.L.T. [20]. Additionally, our method is nearly 2×2$\times$2 × faster speeds for both training and inference compared to W.A.L.T. Under identical training architecture and computational resources, our model processes each training iteration in just $380$ milliseconds, compared to $750$ milliseconds for W.A.L.T. This significant reduction in training time is due to the shorter sequence lengths enabled by our compact factorized representation, making the model both efficient and effective for video generation tasks. Additionally, our factorization and recombination layers are lightweight, adding minimal overhead to the encoding and decoding processes. Additional timing details are included in the appendix.
4.1.3 Training and architecture details
To ensure a fair comparison against W.A.L.T. [20], we adopt a similar training setup and network architecture. The outer layers of our autoencoder architecture mirror the MAGVIT-v2 design to facilitate consistent benchmarking and reliable performance evaluations. In line with recent approaches in latent diffusion models, we simplify the autoencoder by removing the quantizer, working directly with continuous latent representations. To obtain the planes from the latent feature volume, we use average pooling as the factorization operations $\Lambda_{h}$ , $\Lambda_{w}$ and $\Lambda_{t}$. To recompose the planes into feature volume, we use the concatenation operation as $\mathrm{Combine}$. The autoencoder is trained with a combination of objectives, including an L2 reconstruction loss, a perceptual loss, and an adversarial loss, to ensure high-quality latent representations that preserve both fine details and overall structure. The L2 loss encourages accurate pixel-level reconstructions, while the perceptual loss focuses on retaining high-level semantic features, making the representation more robust. Additionally, the adversarial loss sharpens the output, helping the autoencoder produce visually realistic results [15].
To train the diffusion model we use a network architecture identical to W.A.L.T. We use a self-conditioning [8] rate of $0.9$, AdaLN-LoRA [20] with $r=2$ as the conditioning mechanism and zero terminal SNR [33] to avoid mismatch between training and inference arising from non-zero signal-to-noise ratio at the final time in noise schedules. We additionally use query-key normalization in the transformer to stabilize training. Our model is trained with a batch size of $256$ using an Adam optimizer with a base learning rate of $5\times 10^{-4}$ with a linear warmup and cosine decay.
Figure 4: Interpolation results. We show the $7$ interpolated frames for two scenes from the DAVIS-7 [28] dataset, our method generates realistic videos with sharp, detailed frames, achieving quality comparable to VIDIM [28].
4.2 Video interpolation
The video interpolation task evaluates our model’s ability to generate intermediate frames between given keyframes, testing both its temporal coherence and smoothness across sequences. Video interpolation is essential for applications requiring fluid motion reconstruction, such as frame-rate upsampling and video inpainting. In this experiment, we leverage our factorized latent representation to generate the spatio-temporal planes, $\mathbf{P}_{xt}$ and $\mathbf{P}_{yt}$, conditioned on the spatial planes, $\mathbf{P}_{xy}^{1}$ and $\mathbf{P}_{xy}^{2}$. As discussed in Sec. 3.1, the spatial-planes are obtained using a frozen image tokenizer.
The image tokenizer used in our model is trained on WebLI [9]. We train our video tokenizer, with spatial planes obtained from the image tokenizer, on an internal dataset similar to ViDiM, at a resolution of $256\times 256$ with 9 frames. We train the diffusion model on this same internal dataset, specifically curated for video tasks. The training and architectural setup closely follow the details outlined in Section 4.1.3. To evaluate our model’s interpolation capability, we test on the DAVIS-7 and UCF-7 datasets, as proposed in ViDiM [28]. These datasets consist of 400 videos, each containing 9 frames, and feature scenes with significant and often ambiguous motion.
As shown in Section 4.1.1, we summarize the quantitative results and compare them against various prior works using reconstruction-based metrics such as PSNR, SSIM, and LPIPS. Although these metrics are commonly used, they can penalize alternative, yet plausible, interpolations. To address this limitation, we also report FVD on the entire video, providing a more holistic evaluation of interpolation quality. We observe that our method performs comparably to VIDIM on the reconstruction metrics. Notably, unlike VIDIM—a diffusion-based baseline requiring a two-stage process with an initial base model followed by a super-resolution step—our model achieves $256\times 256$ resolution video generation in a single stage, making it both simpler and more efficient. We present qualitative results on two DAVIS scenes in Figure 4 to illustrate the effectiveness of our approach in video interpolation. Our method demonstrates comparable quality to the state-of-the-art VIDIM model while producing noticeably sharper details than other methods. These results emphasize the strength of our factorized representation in preserving fine textures and achieving high-fidelity frame generation in complex video scenes.
4.3 Ablation studies
To validate the effectiveness of our design choices, we perform ablations on the factorization and combine method as well as number of planes.
4.3.1 Factorization
We experiment with two variations of the factorization operation to optimize model performance. In the first variation, we apply mean pooling for $\Lambda_{h}$, $\Lambda_{w}$, and $\Lambda_{t}$, effectively reducing the dimensions while retaining essential features. In the second variation, we use a linear projection to map the feature dimension to $1$ along the targeted axis. When evaluating both models on the UCF class-conditional task Table 3, our results reveal that the mean pooling approach outperforms the linear projection variant, highlighting its effectiveness in balancing simplicity and performance.
[…]
Mean Pooling TML]E3F6FF38 TML]E3F6FF91.13
Linear Projection 50 89.80
Table 3: Ablation: factorization method. We report FVD and Inception scores on the class-conditional task for the UCF-101 [44] dataset, comparing two factorization approaches: mean pooling and linear projection.
4.3.2 Combine
We also experiment with different variations of the combine operation to assess their impact on performance. Specifically, we test two approaches, concatenation and summation:
$\mathbf{V}(x,y,t)$ $=[\mathbf{f}^{1}_{xy}\lvert\lvert\mathbf{f}^{2}_{xy}\lvert\lvert\mathbf{f}_{yt}\lvert\lvert\mathbf{f}^{1}_{xt}]$
$\mathbf{V}(x,y,t)$ $=\mathbf{f}^{1}_{xy}+\mathbf{f}^{2}_{xy}+\mathbf{f}_{yt}+\mathbf{f}^{1}_{xt}$
We evaluate these variations on the UCF class-conditional task and present the results in Table 4. We find that concat yields a better performance as compared to summing the features of from the planes for volume creation. This improvement likely stems from concatenation’s ability to retain more distinct feature information from each plane.
[…]
Concat TML]E3F6FF38 TML]E3F6FF91.13
Sum 45 90.76
Table 4: Ablation: Combine method. We report FVD and Inception scores on the class-conditional task for the UCF-101 [44] dataset comparing performance across two different combination methods.
[…]
Four-plane TML]E3F6FF38 TML]E3F6FF91.13
Tri-plane 52 90.46
Table 5: Ablation: number of planes. We report FVD and Inception scores on the class-conditional task for the UCF-101 [44] dataset comparing performance between tri-plane and four plane representation.
4.3.3 Triplane representations
A common approach to reduce dimensionality in volumetric data is the tri-plane representation [7]. While this representation can be effective for class-conditional generation, it is inherently limited for tasks like frame prediction and video interpolation due to the mixing of temporal information within the xy-plane. PVDM [55] employs this three-plane structure, achieving an FVD of $343$ and an Inception score of 74.40 on the UCF-101 dataset at a resolution of $256\times 256$. Our model achieves a significantly better FVD of $104$ and Inception score of $89.96$ on the same task. PVDM however used a 2D UNet based architecture compared to a transformer diffusion architecture in our model. Thus, to accurately assess the impact of the four-plane representation we perform an experiment where we substitute the four-plane represenation with tri-plane. We report the results in Table 5. Comparatively, the four-plane representation achieves better performance while also providing additional flexibility in its ability to be applied towards frame-conditional tasks in a straightforward manner.
5 Conclusion
In this work, we introduced a factorized latent representation that encodes videos into a four-plane structure, paving the way for more efficient representation of spatiotemporal signals. Coupled with transformer-based diffusion models, our approach enables up to 2×2$\times$2 × speedup in training and inference over models operating directly on volumetric latent features—without compromising performance. Our experiments validate that this representation achieves results on par with the previous state-of-the-art across diverse tasks, including class-conditional generation, video extrapolation and interpolation. This work presents a simple and effective way to improve the efficiency of models that work with volumetric latent spaces.
Acknowledgements
We would like to thank Thomas Kipf for helpful discussions and feedback.
Appendix A Implementation details
A.1 Video autoencoder
To incorporate image data into the training of the video autoencoder, we adopt an image pretraining strategy commonly employed in prior works [51, 53, 20]. Specifically, we first train an image autoencoder using 2D convolutional layers. The trained weights are then used to initialize the video autoencoder. Following the approach in MAGVITv2 [53], which shares a similar architecture with our model, we inflate the 2D weights to 3D by initializing the 3D filters to zero and assigning the last slice of the 3D filter to the corresponding 2D filter weights. This method ensures a smooth transition from image-based training to video-based learning, leveraging the pre-trained image representations effectively.
For class conditional and frame prediction task, we train the tokenizer for $270,000$ iterations with a batch size of $256$. The resulting autoencoder achieves a reconstruction performance of $27.11$ PSNR and $0.829$ SSIM on videos with $128\times 128$ resolution and $17$ frames.
For the video interpolation task, the autoencoder is trained for $450,000$ iterations with the same batch size of $256$. It achieves a reconstruction PSNR of $25.58$ and SSIM of $0.717$ on videos with $256\times 256$ resolution and $9$ temporal frames.
A.2 Denoiser
We use the same transformer architecture across all three tasks, following the design and hyperparameters outlined in W.A.L.T. [20].
  • Class-conditional generation: The denoiser is trained for $74,000$ iterations with a batch size of $256$. The input sequence has a length of $672$, comprising two spatial planes with a resolution of $16\times 16$ each and two spatio-temporal planes with a resolution of $5\times 16$ each.
  • Frame prediction: The denoiser is trained for $270,000$ iterations with a batch size of $256$. The input sequence has a length of $416$, composed of one spatial plane with a resolution of $16\times 16$ and two spatio-temporal planes with a resolution of $5\times 16$ each. The conditioning sequence has a length of $256$, formed by flattening the first spatial plane, which contains information equivalent to the first two latent frames used as conditioning in W.A.L.T.
  • Video interpolation: The model is trained for $100,000$ iterations with a batch size of $256$. The target sequence has a length of $96$, corresponding to the two spatio-temporal planes, while the conditioning sequence has a length of $512$.
A.3 Diffusion
During training, we adopt a scaled linear noise schedule [38] with $\beta_{0}=0.0001$ and $\beta_{T}=0.002$, utilizing a DDPM sampler [22] for the forward diffusion process. During inference, we switch to a DDIM sampler [43] with $50$ steps.
Appendix B Timing details
A detailed timing breakdown for various components of the model during training, measured with different batch sizes on TPU v5e, TPU v4, V100, and A100 devices in the class-conditional generation setting, is provided in Figure 5. These timings were obtained using a model with 214M parameters, alternating between our factorized latent representation and the volumetric latent baseline. For each plot, timings are reported up to the maximum batch size supported on each device. The timings reported in Section 4.1.2 correspond to a model trained on a $4\times 8$ TPU v5e architecture with a batch size of $256$. These measurements approximately align with the timings for a batch size of $8$ shown in row 1 of Figure 5.
Across all devices, our model supports larger batch sizes due to its reduced memory requirements. For instance, on TPU v5e (Figure 5), our model accommodates a batch size of $18$, whereas the baseline is limited to $10$.
The decoder network incurs slightly higher execution time because it contains nearly twice the parameters of the encoder. Although the encoder and decoder in our model are marginally slower than the baseline autoencoder due to the additional factorization and recomposition operations, these operations are executed only once, compared to the denoiser network which is run for $50$ steps during inference, keeping their overall impact minimal.
Figure 5: Timing Breakdown. Execution times for the encoder, denoiser, and decoder are reported across varying batch sizes on TPU architectures (v5e and v4) in Rows 1 and 2, and GPU architectures (V100 and A100) in Rows 3 and 4. The comparison includes timings for factorized latents (blue) and volumetric latents (orange), measured up to the maximum batch size supported without running out of memory (OOM) for each configuration. The timings are reported for a single step measured during training.
Appendix C Qualitative results
We provide frame interpolation results in frameprediction-video folder. Although our frame prediction model achieves a slightly higher FVD, it still generates plausible and visually coherent extrapolations.. We also provide additional results for interpolation on the DAVIS dataset in the interpolation-videos folder.
Appendix D Limitations
Our model was trained on $128\times 128$ resolution videos with 17 frames for the class-conditional generation and frame prediction. Our models performance on longer-duration videos remains unexplored. Similarly, the interpolation model was trained in a single stage. While it achieves competitive results, its quality could potentially be improved by scaling up the dataset, increasing the model size, and incorporating an upsampling stage.